Find the final volume of a gas sample (5 mol) at 1720 mmHg and 64°C of the gas is added, temperature drops to 10 °C and the pressure increases to 1900 mmHg .
1
Expert's answer
2020-06-09T07:21:57-0400


(P1V1) / (T1) = ( P2 V2) / (T2 )



(1720) (5) / (337) = ( 1900) (V2 )/(283)


V2 = ( 283)(1720)(5) / (337)(1900)


V2 = (2433800) / (640300)


V2 = 3.8 mol